### v1

# The Law of Multi-Model Collaboration: Scaling Limits of Model Ensembling for Large Language Models

Abstract
Recent advances in large language models (LLMs) have been largely driven by scaling laws for individual models, which predict performance improvements as model parameters and data volume increase. However, the capabilities of any single LLM are inherently bounded. One solution originates from intricate interactions among multiple LLMs, rendering their collective performance surpass that of any constituent model. Despite the rapid proliferation of multi-model integration techniques such as model routing and post-hoc ensembling, a unifying theoretical framework of performance scaling for multi-model collaboration remains absent. In this work, we propose the Law of Multi-model Collaboration, a scaling law that predicts the performance limits of LLM ensembles based on their aggregated parameter budget. To quantify the intrinsic upper bound of multi-model collaboration, we adopt a method-agnostic formulation and assume an idealized integration oracle where the total cross-entropy loss of each sample is determined by the minimum loss of any model in the model pool. Experimental results reveal that multi-model systems follow a power-law scaling with respect to the total parameter count, exhibiting a more significant improvement trend and a lower theoretical loss floor compared to single model scaling. Moreover, ensembles of heterogeneous model families achieve better performance scaling than those formed within a single model family, indicating that model diversity is a primary driver of collaboration gains. These findings suggest that model collaboration represents a critical axis for extending the intelligence frontier of LLMs.
1 Introduction
In recent years, large language models (LLMs) have made progress under the guidance of a unified theoretical framework: scaling laws that describe performance as a power-law function of model parameters, training data, and computation (Kaplan et al., 2020; Henighan et al., 2020). These laws provide both forecasting power—letting practitioners predict performance before training begins—and a conceptual framework that treats intelligence progress as movement along a resource-constrained performance frontier. Refinements like compute-data trade-offs (Hoffmann et al., 2022) have cemented single-model scaling laws as foundational theory for LLM development. Yet these laws do more than summarize trends: they articulate boundaries of the single model. By characterizing how loss behaves asymptotically as scale increases, they implicitly define what a single model can achieve under ideal conditions (Yang et al., 2025; Grattafiori et al., 2024). But this boundary has theoretical limitations. As models grow, performance gains show diminishing returns, demanding exponentially more resources for marginal improvements. Meanwhile, any single LLM is affected by its random initialization, architecture bias, and training data distribution, resulting in uneven performance across different tasks and domains, even among models of the same size (Bender et al., 2021; Liang et al., 2022). This suggests that single-model scaling laws capture only one dimension along which intelligence can extend.
Motivated by these constraints, various multi-model integration approaches have been proposed (Chen et al., 2025b)—including dynamic routing (Qian et al., 2025), ensemble voting (Zhao et al., 2025), tool orchestration (Zhang et al., 2025), and agent-based collaboration (Tran et al., 2025; Chen et al., 2025a; Li et al., 2024; Fan et al., 2025)—which demonstrate superior performance and efficiency over individual models across diverse application scenarios. But these successes remain empirical and method-specific, justified through heuristics or task evaluations rather than principled understanding of why multi-model collaboration works. The result is a serious mismatch between theory and practice: although the scaling law explains the growth behavior of individual models, there is no similar framework for multi-model collaboration. Does the multi-model collaborative system scale with the total number of parameters in the model pool? Can the performance upper limit of a multi-model collaborative system be predicted based on the total number of parameters in the model pool? Besides the total number of parameters in the model pool, what other factors affect the performance of multi-model collaborative systems? Without theoretical grounding, we cannot compare multi-model systems to large single models rigorously or reason about their ultimate limits. This gap motivates a general law for multi-model collaboration, parallel to existing single-model scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; Kim et al., 2025).
We propose the Law of Multi-model Collaboration, which characterizes the performance limit of multi-model collaborative systems as a function of the total number of parameters of all models involved. Rather than focusing on specific ensemble mechanisms, we adopt a method-agnostic view treating multi-model systems as collective entities. This study operates within the AI Flow framework (An et al., 2025; Shao and Li, 2025), which advocates distributed intelligence integrating artificial intelligence with communication systems through collaborative interactions across heterogeneous models deployed at device, edge, and cloud levels (Yuan et al., 2025). Understanding fundamental performance limits of multi-model collaboration becomes prerequisite for principled system design in this paradigm. Our formulation centers on a model pool of LLMs with total parameter count summed across the pool. For a given input distribution, we characterize the minimal achievable loss through model collaboration, defining a theoretical lower bound analogous to single-model scaling laws. To isolate intrinsic collaboration limits, we consider an idealized oracle integration where per-token cross-entropy loss equals the minimum achieved by any pool model. While unrealizable in practice, this oracle parallels optimality assumptions in prior scaling studies that abstract away algorithmic inefficiencies to reveal fundamental trends. We construct empirical performance frontiers by enumerating model combinations from a fixed pool, evaluating aggregated parameter budgets against achieved loss, and extracting Pareto-optimal points representing the theoretical performance envelope.
Our analysis yields two findings. The multi-model system exhibits a clear power-law scaling between loss and total parameter count, similar to the single model scaling law, but with a steeper downward trend and significantly lower asymptotic loss lower bound, indicating that it utilizes parameters more effectively than the single model scaling law. Additionally, scaling efficiency depends strongly on model diversity: ensembles from different families (Yang et al., 2025; Grattafiori et al., 2024; Team et al., 2025) consistently outperform single-family ensembles under equivalent parameter budgets, suggesting architectural and training heterogeneity drives collaborative gains beyond parameter count. These findings imply model diversity constitutes a fundamental scaling dimension complementary to parameters, data, and compute. The Law of Multi-model Collaboration thus reframes ensemble methods from ad-hoc techniques into a systematic, theoretically grounded pathway for extending artificial intelligence frontiers. This work makes four contributions:
  • We propose the Law of Multi-model Collaboration, the first scaling-law framework characterizing performance limits of LLM ensembles.
  • We demonstrate that multi-model systems follow predictable power-law scaling with respect to total parameter budget, achieving significantly lower theoretical loss bounds than single-model scaling.
  • We show model heterogeneity plays a decisive role in scaling efficiency, with diverse-family ensembles consistently outperforming homogeneous ensembles under equivalent parameter budgets.
  • We establish principled theoretical comparison between single-model and multi-model performance frontiers, reframing ensemble methods as a systematic pathway for extending the intelligence scaling frontier.
2 Problem Formulation and Theoretical Framework
In this section, we formalize the problem setting and introduce the theoretical framework underlying the Law of Multi-model Collaboration. Our goal is to characterize the fundamental performance limits of multi-model systems in a manner that is independent of specific integration algorithms.
2.1 Preliminaries
Let $\mathcal{D}$ denote a fixed text distribution over sequences $x=(x_{1},\dots,x_{T})$. Let $\mathcal{M}=\{M_{1},M_{2},\dots,M_{K}\}$ be a pool of pretrained large language models. Each model $M_{i}$ is associated with a parameter count $P_{i}$ and a tokenizer $\tau_{i}$, which maps an input text $x$ to a token sequence
$\tau_{i}(x)=(t^{(i)}_{1},\dots,t^{(i)}_{n_{i}(x)}),$
where the resulting token sequence length $n_{i}(x)$ may vary across models due to tokenizer differences.
We denote by $p_{M_{i}}(t\mid\cdot)$ the next-token predictive distribution induced by model $M_{i}$ under its tokenizer.
2.2 Text-level Loss under Heterogeneous Tokenization
A key challenge in comparing different language models arises from the fact that distinct models employ different tokenization schemes, resulting in different token sequence lengths for the same input text. To enable a fair comparison across models, we define the modeling loss at the text level, rather than at the per-token average level. Specifically, for a given model $M_{i}$ and input text $x$, we define the loss as the sum of token-level cross-entropy losses:
$\mathcal{L}(M_{i},x)\;=\;\sum_{j=1}^{n_{i}(x)}-\log p_{M_{i}}\!\left(t^{(i)}_{j}\,\middle|\,t^{(i)}_{<j}\right)$. (1)
We emphasize that we deliberately use the summed loss rather than the mean loss, since averaging would implicitly normalize by tokenizer-dependent sequence length and obscure genuine differences in modeling efficiency. The summed formulation ensures that each model is evaluated on its own induced token sequence while preserving comparability at the level of complete texts. The expected loss of model $M_{i}$ over the distribution $\mathcal{D}$ is then given by
$\mathcal{L}(M_{i})\;=\;\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}(M_{i},x)\right].$ (2)
2.3 Oracle Multi-model Collaboration
To isolate the intrinsic upper bound of multi-model collaboration, we consider an idealized oracle ensemble. This oracle abstracts away any practical limitations of model integration and serves solely as a theoretical construct for characterizing achievable performance limits. For a given model pool $\mathcal{S}\subseteq\mathcal{M}$ and input text $x$, the oracle ensemble loss is defined as
$\mathcal{L}_{\mathrm{oracle}}(\mathcal{S},x)\;=\;\min_{M_{i}\in\mathcal{S}}\;\mathcal{L}(M_{i},x).$ (3)
That is, for each individual text, the oracle selects the single model within the pool that yields the lowest text-level loss under its own tokenizer. The expected oracle loss over the distribution $\mathcal{D}$ is
$\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})\;=\;\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}_{\mathrm{oracle}}(\mathcal{S},x)\right].$ (4)
We stress that the oracle ensemble does not correspond to any realizable inference-time system. Instead, it defines a best-case lower bound on loss achievable by any multi-model collaboration mechanism operating over the model pool $\mathcal{S}$. This abstraction is directly analogous to idealized assumptions commonly employed in prior scaling-law analyses.
2.4 Aggregated Parameter Budget
For a model pool $\mathcal{S}$, we define its aggregated parameter budget as the sum of parameters across all constituent models:
$P_{\mathrm{total}}(\mathcal{S})\;=\;\sum_{M_{i}\in\mathcal{S}}P_{i}.$ (5)
This quantity serves as the primary scaling variable for multi-model systems, paralleling the role of parameter count in single-model scaling laws.
2.5 Multi-model Scaling Frontier
Given a fixed model pool $\mathcal{M}$, we consider all possible subsets $\mathcal{S}\subseteq\mathcal{M}$ and evaluate their corresponding pairs
$\left(P_{\mathrm{total}}(\mathcal{S}),\;\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})\right).$
Among these points, we identify the Pareto-optimal frontier, consisting of all subsets for which no other subset achieves both lower loss and lower total parameter count. This frontier represents the theoretical performance envelope of multi-model collaboration under the oracle assumption. The object of study in this work is the functional relationship between $P_{\mathrm{total}}(\mathcal{S})$ and the minimal achievable loss on this frontier.
2.6 Law of Multi-model Collaboration
We are now ready to state the Law of Multi-model Collaboration. Empirically, we observe that the Pareto-optimal oracle loss follows a power-law relationship with respect to the aggregated parameter budget:
$\mathcal{L}_{\mathrm{oracle}}(P)\;\approx\;A\,P^{-\alpha}+\mathcal{L}_{\infty},$ (6)
where $A>0$ is a scaling constant, $\alpha>0$ is the scaling exponent, and $\mathcal{L}_{\infty}$ denotes the asymptotic loss floor.
This law generalizes classical single-model scaling laws to the multi-model setting. Crucially, both the exponent $\alpha$ and the loss floor $\mathcal{L}_{\infty}$ depend on the composition of the model pool, with heterogeneous model pools exhibiting systematically larger exponents and lower asymptotic loss than homogeneous ones.
3 Experimental Protocol for Multi-model Scaling
In this section, we describe the experimental protocol used to empirically characterize the scaling behavior of multi-model collaboration. The design of this protocol closely follows the methodology established in prior single-model scaling law studies, while extending it to the multi-model setting defined in Section 2.
3.1 Model Pool Construction
We construct a model pool $\mathcal{M}$ consisting of $71$ publicly available open-source base language models. The pool spans a wide range of parameter scales and includes multiple widely adopted model families, such as Qwen, LLaMA, Gemma, among others. All models are evaluated in their base (pre-alignment) form to minimize confounding effects introduced by instruction tuning or reinforcement learning from human feedback. The selected models exhibit substantial diversity in architecture, training data composition, and optimization strategies, resulting in a broad spectrum of modeling capabilities. This diversity enables a systematic investigation of both homogeneous and heterogeneous multi-model collaboration.
3.2 Evaluation Dataset and Loss Measurement
All models are evaluated on a fixed held-out text dataset sampled from the distribution $\mathcal{D}$. For each model $M_{i}\in\mathcal{M}$ and each input text $x$, we compute the text-level loss $\mathcal{L}(M_{i},x)$ as defined in Equation (1), namely the sum of token-level cross-entropy losses under the model’s own tokenizer. The expected loss $\mathcal{L}(M_{i})$ is estimated by averaging $\mathcal{L}(M_{i},x)$ over the evaluation dataset. This procedure yields a single-model performance baseline that is directly comparable across models despite heterogeneous tokenization schemes.
3.3 Enumeration of Multi-model Combinations
To explore the space of multi-model collaboration, we enumerate all unordered pairwise combinations of models from the pool $\mathcal{M}$. For each pair $\mathcal{S}=\{M_{i},M_{j}\}$, we compute:
  • the aggregated parameter budget $P_{\mathrm{total}}(\mathcal{S})=P_{i}+P_{j}$, and
  • the oracle ensemble loss $\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})$ as defined in Equation (3).
Each model pair is further categorized into one of two groups:
  1. 1.
    Homogeneous combinations, where both models belong to the same model family;
  2. 2.
    Heterogeneous combinations, where the models originate from different model families.
This partition enables a controlled comparison of scaling efficiency with and without architectural and training diversity.
3.4 Pareto-optimal Frontier Extraction
Due to the large number of models and the wide variance in their capabilities, the raw set of evaluated points contains many dominated configurations. To focus on the intrinsic performance limits, we extract Pareto-optimal frontiers separately for:
  • single-model configurations ($\mathcal{S}=\{M_{i}\}$),
  • homogeneous multi-model combinations, and
  • heterogeneous multi-model combinations.
A configuration $\mathcal{S}$ is considered Pareto-optimal if no other configuration achieves both a lower oracle loss and a lower aggregated parameter budget. The resulting Pareto frontiers represent the empirical lower envelopes of achievable loss under the oracle collaboration assumption.
Figure 1: Comparison between scaling performance under two settings: single model and multi-model ensemble. Solid curves denote nonlinear least-squares fits of the scaling formula, while points correspond to Pareto-optimal lower envelopes. Model ensembling requires fewer parameters to achieve a target loss, and reduces the lower bound of the test loss significantly.
3.5 Scaling Law Fitting
For each Pareto-optimal frontier, we fit the scaling law of the form
$\mathcal{L}(P)=A\,P^{-\alpha}+\mathcal{L}_{\infty},$ (7)
as introduced in Equation (6). The parameters $(A,\alpha,\mathcal{L}_{\infty})$ are estimated using nonlinear least squares regression on the Pareto-optimal points. Fitting is performed in the original loss domain rather than the log-transformed domain to ensure consistency with prior scaling-law analyses. This procedure yields comparable scaling exponents and asymptotic loss estimates for single-model, homogeneous multi-model, and heterogeneous multi-model settings, enabling a direct empirical comparison of their scaling behaviors.
4 Empirical Scaling Laws of Multi-model Collaboration
In this section, we present empirical evidence supporting the Law of Multi-model Collaboration. By analyzing the Pareto-optimal performance frontiers constructed in Section 3, we demonstrate that multi-model systems exhibit stable and predictable power-law scaling behavior with respect to aggregated parameter budgets, and that this behavior differs systematically from that of single models.
4.1 Scaling Behavior of Single and Multi-model Systems
Figure 1 and Figure 2 illustrates the Pareto-optimal lower envelopes for three configurations: single models, homogeneous multi-model ensembles composed of models from the same family, and heterogeneous multi-model ensembles composed of models from different families. Each point on the envelope corresponds to a Pareto-optimal configuration in terms of aggregated parameter count and oracle loss, while the solid curves represent nonlinear least-squares fits of the scaling law defined in Equation (6). Across all three settings, we observe a clear power-law relationship between loss and total parameter count. This result indicates that, despite increased structural complexity, multi-model systems exhibit regular and predictable scaling behavior analogous to classical single-model scaling laws.
Table 1: Parameter fitting results under different model collaboration settings. The loss floor $L_{\infty}$ is significantly reduced for model ensembles across families, signifying the benefit of multi-model collaboration.
Collaboration settings Raw points Pareto points $L_{\infty}$ AA $\alpha$
Single model 71 12 1765.0285 886.1545 0.3578
Model ensemble within a family 506 22 1780.0938 979.6996 0.8500
Model ensemble across families 1979 32 1625.5834 948.0984 0.5516
4.2 Quantitative Comparison of Scaling Exponents
To quantify differences in scaling efficiency, we fit the scaling law to the Pareto-optimal frontiers in each configuration. Table 1 reports the fitted parameters, including the scaling exponent $\alpha$ and the asymptotic loss floor $L_{\infty}$. The Single-model scaling yields an exponent of $\alpha=0.3578$, consistent with previously reported scaling behavior in language modeling. In contrast, multi-model systems achieve substantially larger scaling exponents. Heterogeneous ensembles composed of models from different families attain an exponent of $\alpha=0.5516$, indicating a significantly faster rate of loss reduction per unit increase in total parameters. Homogeneous ensembles within the same model family exhibit an even steeper exponent of $\alpha=0.8500$, reflecting rapid initial gains at smaller aggregated parameter budgets. These results demonstrate that multi-model collaboration enables more efficient utilization of parameters than single-model scaling, as evidenced by the consistently larger scaling exponents. Specifically, the fitted scaling laws corresponding to the three collaboration settings are given by:
$\mathcal{L}_{\text{single}}(P)$ $=\frac{886.1545}{P^{0.3578}}+1765.0285,$ (8)
$\mathcal{L}_{\text{same}}(P)$ $=\frac{979.6996}{P^{0.8500}}+1780.0938,$ (9)
$\mathcal{L}_{\text{diff}}(P)$ $=\frac{948.0984}{P^{0.5516}}+1625.5834,$ (10)
where $P$ denotes the total parameter count, and the subscripts indicate single-model, same-family ensemble, and cross-family ensemble settings, respectively.
4.3 Asymptotic Loss Floors
Beyond scaling speed, the fitted asymptotic loss floors $L_{\infty}$ reveal important qualitative differences between the three regimes. Single-model scaling exhibits an estimated loss floor of $L_{\infty}=1765.03$. Homogeneous multi-model ensembles yield a comparable loss floor of $L_{\infty}=1780.09$, suggesting limited asymptotic improvement when combining models within the same family. In contrast, heterogeneous ensembles achieve a substantially lower asymptotic loss floor of $L_{\infty}=1625.58$. This pronounced reduction indicates that cross-family collaboration enables access to complementary modeling capabilities that are unattainable by scaling a single architecture or combining closely related models.
4.4 Interpretation of Homogeneous versus Heterogeneous Scaling
Although homogeneous ensembles exhibit a steeper scaling exponent, their advantage is primarily confined to smaller parameter regimes. As the aggregated parameter budget increases, their performance rapidly saturates at a loss floor comparable to that of single models. This behavior suggests that homogeneous ensembles primarily reduce redundancy within a shared inductive bias, but do not fundamentally extend modeling capacity. Heterogeneous ensembles, by contrast, achieve both a faster scaling rate than single models and a significantly lower asymptotic loss floor. This combination indicates that architectural and training diversity enables persistent gains across scales, effectively shifting the performance frontier downward rather than merely accelerating convergence toward an existing limit.
4.5 Evidence for the Law of Multi-model Collaboration
Taken together, these results constitute empirical evidence for the Law of Multi-model Collaboration. Multi-model systems exhibit predictable power-law scaling with respect to aggregated parameter budgets, while achieving systematically improved scaling efficiency and lower theoretical loss bounds compared to single-model scaling. Crucially, the composition of the model pool plays a central role in determining scaling behavior. Diversity across model families emerges as a key factor in extending the attainable performance frontier, highlighting multi-model collaboration as a complementary and orthogonal scaling axis to parameters, data, and compute.
Figure 2: Comparison between scaling performance between model ensemble within the same family and that across different families. Solid curves denote nonlinear least-squares fits of the scaling formula, while points correspond to Pareto-optimal lower envelopes. Collaboration beyond a single model series results in better scaling performance, validating the complementarity among different model series.
5 Interpretation and Theoretical Implications
In this section, we interpret the empirical results given in Section 4 and discuss their broader theoretical implications. Instead of proposing new algorithms, we want to know what the observed scaling behavior can tell us about how different models work together and if it matches up with any ideas people already have about how models should behave when they get bigger.
5.1 Multi-model Collaboration as an Extension of Scaling Laws
Classical scaling laws characterize how the performance of a single model improves as a function of its parameter count, data, and compute. Implicit in these laws is the assumption that scaling occurs along a single architectural and representational trajectory (Kaplan et al., 2020; Hoffmann et al., 2022). Our results indicate that multi-model collaboration is a qualitatively distinct kind of scaling.There is a stable power law relationship between oracle loss and aggregated parameter budget for multi-model system which means that multi-model system has its own scaling regime. This regime is not just a simple pile-up of single-model behavior. On the contrary, better scaling exponents and lower asymptotic loss floors show that working together lets us reach parts of the performance frontier that we can’t get to just by making one model bigger. And so this leads us to think of the total parameter budget of a pool of models as a new kind of scaling variable that’s somewhat similar to the number of parameters for a single model, but it’s living in a much bigger model space.
5.2 Interpretation of Scaling Exponents
The scaling exponent $\alpha$ reflects the efficiency with which additional parameters translate into performance gains. The consistently larger exponents observed for multi-model systems indicate that parameters distributed across multiple models can be utilized more efficiently than parameters concentrated within a single model. Notably, homogeneous multi-model ensembles show extremely steep scaling exponents for small parameter budgets, indicating quick early improvements.But it has early saturation too, with asymptotic loss floors that match those of a single model.This pattern implies that merging similar models mainly reduces redundancy or variance within a common inductive bias, rather than substantially enlarging representational capacity.In contrast, heterogeneous ensembles get better scaling exponents and much smaller asymptotic loss floors.This combination implies that architectural and training diversity enables sustained gains across scales, consistent with the notion that different models encode complementary hypotheses about the data distribution (Dietterich, 2000).
5.3 Lower Loss Floors and the Limits of Single-Model Scaling
The reduction of the asymptotic loss floor that we see with heterogeneous multi-model systems is significant for our understanding of the limits on scaling single models.In classical scaling laws, the loss floor is often interpreted as reflecting irreducible error under a fixed model class and training regime (Kaplan et al., 2020; Henighan et al., 2020).From our findings we can conclude that it’s not just the number of parameters which determines this floor, but also the variety of inductive biases at the disposal of the system. By allowing multiple model families to contribute, multi-model collaboration effectively expands the hypothesis space, thereby lowering the attainable error bound (Domingos, 2012). So from this point of view, the loss floor for single-model scaling laws needs to be understood as being tied to some particular lineage of architecture, rather than as some sort of hard stop.
5.4 Relation to Existing Paradigms
Multi-model collaboration shares superficial similarities with ensemble learning and mixture-of-experts methods, but differs in important respects. Traditional ensembles (Chen et al., 2025b) are typically designed to reduce variance or improve robustness around a fixed model class, while mixture-of-experts (Cai et al., 2025) architectures are trained jointly with explicit routing mechanisms. In contrast, the framework studied here considers pretrained models as fixed entities and focuses on the theoretical limits of their collective performance under idealized coordination.The oracle formulation removes the training dynamics and routing mechanism from consideration, so that the scaling behavior itself can be studied on its own.Therefore, the Law of Multi-model Collaboration can be regarded as supplementary to, but not a substitute for, existing single-model scaling theories.
5.5 Implications for the Intelligence Frontier
Taken together, these results suggest that progress in language modeling need not rely exclusively on scaling individual models along a single dimension.Model diversity becomes a new orthogonal direction that the intelligence frontier can expand into. The Law of Multi-Model Collaboration transforms multi-model systems from heuristic engineering constructions into objects that follow consistent, predictable scaling rules. It means that we could make progress in the future without necessarily having to train bigger models, but rather by cleverly combining different kinds of models so they can work well together.
6 Limitations and Scope of the Law
While the empirical results presented in this work provide strong evidence for the Law of Multi-model Collaboration, it is important to clarify the scope of the law and acknowledge its limitations.These limitations arise from the theoretical perspective taken here, and they do not detract from the main contribution, which is identifying the basic performance limits rather than to suggest practical systems.
6.1 Idealized Oracle Assumption
An important abstraction in our formulation is the oracle ensemble, which selects, for each input text, the model within a pool that achieves the lowest loss.It’s an intentionally idealized assumption that does not correspond to any feasible inference time mechanism.In practice, it is usually impossible to determine which model performs best for each input without knowing the ground-truth loss.However, this abstraction serves a role analogous to idealized assumptions commonly employed in prior scaling-law analyses, such as optimal data usage or perfect optimization.The oracle ensemble sets an upper limit on the performance that can be achieved with multi-model collaboration, thus enabling the study of intrinsic scaling behavior without considering algorithmic inefficiency. Therefore our results should be seen as theoretical limits rather than deployable architectures.
6.2 Upper Bounds versus Realizable Systems
The Law of Multi-model Collaboration describes the behavior of Pareto-optimal lower envelopes under an oracle assumption. As such, it does not claim that practical multi-model systems can attain the same performance. Real world implementations have other constraints such as routing errors, coordination overhead, latency, and compute costs.Nonetheless, upper-bound analyses play a critical role in guiding system design by revealing what is possible in principle. Like how single-model scaling laws give us ideas on what kind of models to use and where to put the data even though they make some idealized assumptions, this law gives us something to compare against and encourage people to work on better ways of putting different models together.
6.3 Restriction to Pairwise Model Combinations
Our empirical analysis focuses on pairwise combinations of models. This restriction is due to computational constraints; the number of potential subsets increases exponentially as the size of the model pool grows.Higher order combinations may give even better results, but the scaling behavior is already evident for pairs. Importantly, the law itself is not inherently limited to pairs. The formulation in Section 2 applies to arbitrary model subsets, and extending the empirical analysis to larger ensembles is a natural direction for future work.Higher-order combinations are expected to make the observed trends more pronounced, possibly resulting in even lower performance frontiers.
6.4 Dependence on the Model Pool
The fitted scaling parameters rely on the makeup of the underlying model pool.Different choices for model families, parameter ranges, or training protocols can cause differences in scaling exponents and loss floors.Accordingly, the numerical values reported in this work should not be interpreted as universal constants.Rather, the law asserts the existence of a stable power-law relationship and systematic differences between single-model and multi-model scaling regimes. The qualitative conclusions about enhanced scaling efficiency and decreased asymptotic loss with heterogenous collaboration hold up well for all kinds of models.
6.5 Loss Definition and Tokenization Effects
To make it possible for different models with different tokenizers to have a fair comparison, we evaluate the text-level loss by adding up all the token-level cross entropy values instead of taking their average. This choice avoids the normalization artifacts from tokenizer-dependent sequence lengths, but it means that the loss values depend on how we tokenize things.While this formulation is well-suited for comparing oracle performance bounds, alternative loss normalizations may be more appropriate in other contexts. The law itself is independent from the particular choice of the loss function, and it will be interesting for future work to study the effect of different evaluation metrics on the observed scaling behavior.
6.6 Scope of Applicability
Lastly, we stress that the Law of Multi-model Collaboration is meant to supplement, not supplant, current single-model scaling theories.Single model scaling still matters for optimization, training dynamics, and resource allocation inside one fixed architecture.The law proposed here operates at a different level of abstraction, focusing on how multiple pretrained models interact as collective entities.Within this scope, the law provides a principled framework for reasoning about multi-model systems and highlights collaboration and diversity as fundamental dimensions of scaling.
7 Conclusion
We introduce the Law of Multi-model Collaboration, a scaling law defining the theoretical performance limits of systems composed of multiple pretrained LLMs. By employing an oracle-based, method-agnostic formulation, we isolate the intrinsic behavior of multi-model collaboration and provide a principled comparison to classical single-model scaling laws. Our empirical study across diverse open-source models demonstrates that multi-model systems follow stable power-law scaling relative to aggregated parameter budgets. We have conducted a systematic empirical study on various open-source language models, which shows that multi-model systems show stable power-law scaling regarding their combined parameter budget. Compared to single-model scaling, multi-model collaboration can achieve better scaling efficiency, and in heterogenous pools, it has much lower asymptotic loss floors.These results suggest that distributing parameters across multiple models—especially those with diverse inductive biases—can be more effective than monolithic scaling within a single architecture. Beyond the empirical results, this work reframes multi-model systems as objects governed by regular and predictable scaling behavior rather than as ad-hoc engineering solutions.The proposed law identifies model diversity as a fundamental scaling dimension, alongside parameters, data, and compute. From this perspective, the perceived limits of single-model scaling appear to be conditional on architectural lineage rather than absolute boundaries. While the Law of Multi-model Collaboration defines theoretical upper bounds rather than realizable systems, such bounds are essential for clarifying what is achievable in principle. We hope this work inspires further research into diversity-aware scaling, principled integration mechanisms, and broader theories of collective intelligence in machine learning.